Blocks A and B are moving toward each other. A has a mass of 6.7

and a velocity of 6.5m/s, while B has a mass of 9.2kg and a velocity of −5.9m/s. They suffer a completely inelastic collision. During the collision, how much kinetic energy is lost? Only give the numerical value and don't include the unit


1
Expert's answer
2021-12-12T16:41:12-0500
"(6.7+9.2)V=6.7(6.5)-9.2(5.9)\\V=-0.67\frac{m}{s}\\\Delta E=0.5(6.7(6.5)^2)+0.5(9.2(5.9)^2)\\-0.5((6.7+9.2)0.67^2)=300\ J"